OpenStudy (firejay5):

The sum of 3 number is 20. The second number is 4 times the first, and the sum of the first and third is 8. Find the numbers.

OpenStudy (anonymous):

a+b+c=20 b=4a a+c=8 The first equation minus the second equation gives you b=12 Plugging that into the second equation then gives you a=3 Plugging into the third equation gives you c=5

OpenStudy (firejay5):

like your steps

OpenStudy (anonymous):

i meant first minus third in the first step. but thanks :)

OpenStudy (firejay5):

can you show me your steps though @timeparadox

OpenStudy (anonymous):

like how i came up with the equations, or after that?

OpenStudy (firejay5):

like your explaining and your steps on how you got your answer

OpenStudy (anonymous):

ok i'll try to go into more detail: assume that your three numbers are a, b, and c. they add up to 20; therefore, a+b+c=20. second number is 4 times the first; therefore, b=4*a first plus third is 8; therefore, a+c=8 now you can subtract the first and third equations:  plugging that into the second equation for b, 12=4*a. therefore, 12/4=a=3 plugging that into the third equation means 3+c=8 c=8-3=5 hope that helped!